# SpotLight: Shadow-Guided Object Relighting via Diffusion

### v2

@@ Abstract
Recent work has shown that diffusion models can serve as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise lighting control can be achieved for object relighting simply by providing a coarse shadow of the object. Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, leverages existing neural rendering approaches and achieves controllable relighting results with no additional training. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting. We also demonstrate other applications, such as hand-scribbling shadows and full-image relighting, demonstrating its versatility.
@@ 1 Introduction
Image relighting—the process of modifying the lighting conditions of a scene—is a capability that has long been confined solely to the field of computer graphics, where a user is free to control the lighting conditions in the 3D scene before launching the rendering process. When one is given a photograph, however, no such control is possible. Relighting then becomes a hard task because of the complex interactions between light, material properties, and surface characteristics, where different materials respond distinctly to lighting variations through specular highlights, surface orientation, and texture. The recent rise of powerful image generation models (e.g., [54]) has enabled exciting possibilities for image relighting [31, 76].
This paper focuses on providing lighting control to recent diffusion image relighting frameworks, where one wishes to modify the appearance of a single object. This is particularly useful in the context of object compositing, that is the task of inserting a new (2D or 3D) object into a background image. In this context, achieving control over the relighting results is challenging since the object must be relit in a way that abides by the desired lighting properties (e.g., direction, intensity, etc.), yet it must also realistically harmonize with the new background image.
@@
We present SpotLight, a method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [74, 29]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer. Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it to relight the object accordingly. We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object as it moves around it.
@@
Fig. 1 shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [19] according to the desired lighting direction. From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting. To achieve this, SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object.
Our core technical contribution lies in our shadow guidance approach, which refines the guiding shadow input while realistically relighting the virtual object. Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control. This builds upon existing object compositing networks [78, 75] that operate on intrinsic images along with diffusion models’ learned priors to render realistic outputs. We demonstrate that SpotLight’s guidance mechanism handles complex scenarios and maintains physical plausibility without requiring manual intervention nor any additional training.
@@
In summary, our contributions are as follows. First, we introduce a framework that enhances off-the-shelf diffusion models with object relighting control without requiring additional training. We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models. Finally, we release an evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
@@ 2 Related work
Conditional image generation Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [59, 27, 60, 54]. These models offer control mechanisms for various modalities, such as text [50, 54], image content [55, 21], intrinsic images [78, 43, 75, 39], classifier guidance [16] and cross-modal data [12, 72]. Notably, training-free guidance methods [44, 20, 3, 4, 26] have enabled image editing capabilities without specific training, forming the basis of our approach.
@@
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for harmonization [53, 48, 34, 62] do not prioritize physically accurate shading of the objects. Recently, [49] utilized a single-view multi-illumination dataset [45] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [67, 11] for relighting indoor scenes. IC-Light [76] enforces consistency in appearance to relight portraits and various objects. LumiNet [68] and ScribbleLight [14] achieve image relighting through illumination transfer or user-defined scribbles, respectively. Some methods specialize in outdoor scenes, using geometric priors like depth [25] or normalized coordinates [31] or both [40]. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [9] to harmonization of foreground-background albedo [13]. Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [78, 75, 39], which can be leveraged to achieve zero-shot object composition. However, these methods have limited lighting control. ZeroComp [78] provides no lighting control, RGBX [75] can only provide lighting control through a text prompt, and DiffusionRenderer [39] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set. Other diffusion-based approaches, such as Neural Gaffer [29], DiLightNet [74] and IllumiNeRF [79] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
@@
Explicit lighting estimation approaches infer HDR lighting from a single image [22, 24, 36, 80, 47, 38], though they generally lack controllability. Parametric models [23, 65, 15] or GAN inversion methods [64] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
@@
Shadow generation Generative models for rendering realistic shadows [77, 41, 28, 42], or harmonizing rendered shadows with the image [63] have been proposed. Other methods enable controllable shadows for 3D [56] and 2D objects [57, 58]. Our approach can leverage existing shadow generation methods (e.g., [58], see supp.) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
@@
Image intrinsics Decomposing images into albedo and shading has long been studied [7, 33, 46, 66, 13]. Numerous works have also focused on estimating scene geometry, for example, depth [51, 8, 73, 30, 69, 70] and normals [18, 6, 52, 5, 71]. Recently, approaches such as [36, 80] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [10, 17, 30, 71, 2, 43]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
@@ 3 Background: diffusion renderers
Recent work [75, 78] has shown that diffusion models can be adapted to generate images from intrinsic maps including materials (e.g., albedo, roughness, metallic), geometry (e.g., surface normals and depth), and shading. This capability was built using a ControlNet with ZeroComp [78] or full finetuning with RGBX [75]. We refer to this new class of conditional generation models as “diffusion renderers”.
In these models, the denoising network is ran at each denoising step to obtain
where represents the network -prediction, the noisy latents, and the conditioning intrinsic maps.
It has been shown that these methods can realistically composite virtual objects into images in a zero-shot fashion, despite not being trained for this task. Provided a mask of the virtual object and its intrinsic maps (rendered with graphics shaders), the object can be trivially composited in most intrinsic maps of the background image (estimated with trained estimators). Since computing the virtual object shading requires complex lighting simulations, diffusion approaches estimate the shading by providing a partial shading map [78], or train inpainting models [75]. However, unlike traditional rendering, these methods offer no control over the lighting conditions.
In this work, we use a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, analogous to a spotlight moving around the object. We now describe our approach in detail.
@@ 4 SpotLight
We employ a user-provided, guiding shadow to steer the pre-trained diffusion process, thereby enforcing matching illumination on the inserted object. Our key insight is that even an approximate shadow encapsulates important lighting information. By integrating this cue early in the diffusion process, we simultaneously refine both the composite appearance and its shadow, yielding a natural blend with the scene.
@@ 4.1 Approach overview
Our approach requires two inputs: an object mask and a user-specified guiding shadow . This shadow, computed via fast rasterization or specified through hand-drawn scribbles, is used to guide the latent diffusion process. Without retraining the pre-trained renderer , we steer its generative process by incorporating the guiding shadow into the latent space.
4.2 Blending shadows
To fuse the guiding shadow with the latent representation, we update the noisy latent code at each diffusion timestep, similar to [4]. In our framework, the updated latent
serves as a soft constraint that steers the process toward a shadow-consistent output. Here, is the composite of the object albedo, the target shadow, and the background, and controls the strength of the guiding shadow. The latent shadow mask is obtained by downsampling the target shadow to the latent resolution (see supp. for details).
4.3 Shading the object via dual-branch guidance
A key element of our approach is to amplify the impact of the guiding shadow on the object shading through classifier-free guidance (CFG) [26]. We process two parallel diffusion branches: a positive branch that incorporates the desired shadow direction, and a negative branch that suppresses an undesired shadow direction (e.g., via an opposing light direction). Fig. 2 shows our dual-branch guidance mechanism. Classifier-free guidance then combines these branches:
where sets the guidance scale and is the object mask downsampled to the latent resolution. This dual-branch strategy reinforces the intended lighting on the object while naturally blending it with the background.
@@ 4.4 Final image synthesis
After the guided diffusion process, the resulting latent is decoded with the VAE decoder. The background preservation strategy of [78] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
@@ 5 Evaluation on 3D object compositing
Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene. To address this, we build over the evaluation setup of [78] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas. We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render. In this section, we describe how these scenarios are used for quantitative evaluation and a user study to assess the realism and controllability of the methods.
@@ 5.1 Evaluation dataset
We create datasets for the two scenarios using the Blender scene information graciously provided by the authors of [78]. We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [70]. We also improve the ground truth intrinsic maps by providing correct antialiased versions.
“Reference-based” scenario To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [78] to ensure lighting fidelity around the inserted object. We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [19]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3, in the “ground truth” column. This dataset is intended for quantitative evaluation (sec. 5.4).
“User-controlled” scenario This dataset simulates a scenario where a user specifies a desired light position. 8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images. This dataset is utilized in the user study (sec. 5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.

### v3

@@ Abstract
Recent work has shown that diffusion models can serve as powerful neural rendering engines that can be leveraged for inserting virtual objects into images. However, unlike typical physics-based renderers, these neural rendering engines are limited by the lack of manual control over the lighting, which is often essential for improving or personalizing the desired image outcome. In this paper, we show that precise and controllable lighting can be achieved without any additional training, simply by supplying a coarse shadow hint for the object. Indeed, we show that injecting only the desired shadow of the object into a pre-trained diffusion-based neural renderer enables it to accurately shade the object according to the desired light position, while properly harmonizing the object (and its shadow) within the target background image. Our method, SpotLight, is entirely training-free and leverages existing neural rendering approaches to achieve controllable relighting. We show that SpotLight achieves superior object compositing results, both quantitatively and perceptually, as confirmed by a user study, outperforming existing diffusion-based models specifically designed for relighting. We also demonstrate other applications, such as hand-scribbling shadows and full-image relighting, demonstrating its versatility.
@@ 1 Introduction
Object relighting—the process of inserting a virtual object in an image and lighting it appropriately—has traditionally been confined to 3D graphics pipelines, where lighting is explicitly modeled and rendered. In contrast, object relighting in real photographs remains a fundamentally ill-posed problem since it requires inferring the physical properties of the scene: its geometry, materials, and illumination are not observed, and their complex interactions—such as specular reflections, soft shadows, or material-dependent reflectance—must be inferred or approximated.
Recent advances in image generative models, especially diffusion-based methods [rombach2022high], have opened up new possibilities for object relighting [zhang2025iclight, zhang2025zerocomp, liang2025diffusionrenderer]. These models encode strong priors over the appearance of scenes under varied lighting, allowing implicit reasoning about illumination, material, and geometry. Yet, precise and localized relighting control remains limited, particularly in the context of virtual object insertion, where a single object must be relit.
We focus on the problem of local lighting control for object compositing using generative models. Specifically, we aim to relight an inserted object such that the object conforms to a specified lighting direction and intensity around it, while blending it naturally with the background. Achieving such control is challenging: the relighting must align with physical cues like shadows and shading, yet remain flexible enough to work across varied materials, perspectives, and image contexts. Diffusion models offer powerful priors, but they are not inherently controllable without carefully designed guidance.
@@
We present SpotLight, a training-free method for plausible, realistic, and controllable object relighting by leveraging the rich prior knowledge encoded in pre-trained, diffusion-based neural renderer. Rather than obtaining specialized high-quality data and training a model specifically for the task of object relighting (e.g., [zeng2024dilightnet, jin2024neural_gaffer]), our key insight is to drive the relighting process by supplying a coarse representation of the desired shadow of the object to a general-purpose neural renderer. Not only is a 2D shadow an intuitive input for human control, but we also find that diffusion models can effectively utilize it. We call our method SpotLight because it guides the diffusion process in a manner analogous to a physical spotlight, selectively illuminating the object to insert in the image.
SpotLight exploits classifier-free guidance within existing pre-trained diffusion-based renderers and progressively harmonizes imperfect shadows with the background while relighting the object. Our guidance mechanism steers the rendering process and modifies the appearance of the inserted composite, resulting in local lighting control. This mechanism is applied to recent diffusion models pre-trained to render images from input intrinsic maps (depth, normals, albedo, etc.) [zhang2025zerocomp, zeng2024rgb].
@@
SpotLight: Shadow-Guided Object Relighting via Diffusion shows an example of such shadow-based lighting control: the shadow of the chair is rendered using basic shadow mapping [eisemann2011realtime] according to the desired lighting direction. From this, SpotLight realistically blends the input guiding shadow within the background, and relights the virtual object to match the desired lighting.
@@
In summary, our contributions are as follows. First, we introduce a framework that enhances pre-trained diffusion renderers with object relighting control without requiring additional training. We demonstrate superior quality, both perceptually through user studies and quantitatively, compared to competitive specialized models. Finally, we release an evaluation dataset tailored for lighting control in object insertion, which will be made available upon publication.
@@ 2 Related work
Conditional image generation. Diffusion-based generative models have become the de facto choice for image generation due to their ability to produce high-quality images [sohl2015deep, ho2020denoising, song2020denoising, rombach2022high]. These models offer control mechanisms for various modalities, such as text [ramesh2022hierarchical, rombach2022high], image content [saharia2022palette, gal2022image], intrinsic images [zhang2025zerocomp, luo2024intrinsicdiffusion, zeng2024rgb, liang2025diffusionrenderer], classifier guidance [dhariwal2021diffusion] and cross-modal data [brooks2023instructpix2pix, ye2023ip]. Notably, training-free guidance methods [meng2022sdedit, epstein2023diffusion, avrahami2022blended, avrahami2023blendedlatent, ho2021classifierfree] have enabled image editing capabilities without specific training, forming the basis of our approach.
@@
Image relighting modifies the global or local shading of an image without changing other properties such as geometry and materials. Previous methods for harmonization [reinhard2001color, pitie2005n, lalonde2007using, tsai2017deep] do not prioritize physically accurate shading of the objects. Recently, [poirier2024diffusion] utilized a single-view multi-illumination dataset [murmann2019multi] to enable direct control over the dominant lighting direction. Retinex theory has been leveraged [xing2024retinex, bhattad2024stylitgan] for relighting indoor scenes. IC-Light [zhang2025iclight] enforces consistency in appearance to relight portraits and various objects. LumiNet [xing2025luminet] and ScribbleLight [choi2025scribblelight] achieve image relighting through illumination transfer or user-defined scribbles, respectively. Some methods specialize in outdoor scenes, using geometric priors like depth [griffiths2022outcast] or normalized coordinates [kocsis2024lightit] or both [lin2025urbanir]. These methods however cannot handle lighting on specific region or objects in the scene. Recent work on object relighting has evolved from consistency-based approaches in intrinsic images [bhattad2022cut] to harmonization of foreground-background albedo [careaga2023intrinsic]. Recently, diffusion models have been adapted for the task of rendering from partial intrinsic maps [zhang2025zerocomp, zeng2024rgb, liang2025diffusionrenderer], which can be leveraged to achieve zero-shot object composition. However, these methods have limited lighting control. ZeroComp [zhang2025zerocomp] provides no lighting control, RGBX [zeng2024rgb] can only provide lighting control through a text prompt, and DiffusionRenderer [liang2025diffusionrenderer] requires a prohibitively long and approximate lighting estimation process to annotate each sample in the training set. Other diffusion-based approaches, such as Neural Gaffer [jin2024neural_gaffer], DiLightNet [zeng2024dilightnet] and IllumiNeRF [zhao2024illuminerf] utilize HDR environment maps and multi-lighting renders, but are not designed for object composition in existing images. LightLab [magar2025lightlab] achieves fine-grained control over existing light sources, but does not handle object insertion. Careaga et al. [careagaRelighting] propose a sophisticated inverse rendering method for data generation, wheras SpotLight is training-free. IntrinsicEdit [lyu2025intrinsicedit] proposes a training-free method for lighting-based editing, but, unlike our approach, lacks explicit control over object-specific virtual lighting effects. We bridge this gap by introducing a shadow-guided strategy for object relighting that leverages existing models without any additional training.
@@
Explicit lighting estimation approaches infer HDR lighting from a single image [gardner2017sigasia, garon2019cvpr, li2020cvpr, zhu2022irisformer, phongthawee2024diffusionlight, liang2024photorealistic], though they generally lack controllability. Parametric models [gardner2019iccv, weber2022editable, dastjerdi2023everlight] or GAN inversion methods [wang2022stylelight] offer more control, but these methods require physics-based rendering engines for generating the composite image—here, we focus instead on neural renderers.
@@
Shadow generation. Generative models for rendering realistic shadows [zhang2019shadowgan, liu2020arshadowgan, hong2022shadow, liu2024shadow], or harmonizing rendered shadows with the image [valenca2023shadow] have been proposed. Other methods enable controllable shadows for 3D [sheng2021ssn] and 2D objects [sheng2022controllable, sheng2023pixht]. Our approach can leverage existing shadow generation methods (e.g., [tasar2024controllable]) and shows that diffusion-based neural renderers can be conditioned on such shadows to achieve object relighting.
@@
Image intrinsics. Decomposing images into albedo and shading has long been studied [barrow1978recovering, kovacs2017shading, philip2019multi, wu2023measured, careaga2023intrinsic]. Numerous works have also focused on estimating scene geometry, for example, depth [ranftl2020towards, bhat2023zoedepth, yin2023metric3d, ke2024repurposing, yang2024depth, depth_anything_v2] and normals [eftekhar2021omnidata, bae2021estimating, ranftl2021vision, bae2024rethinking, ye2024stablenormal]. Recently, approaches such as [li2020cvpr, zhu2022irisformer] jointly infer different intrinsic maps including shape, spatially-varying lighting, scene geometries, and materials. Conversely, generative models have shown a powerful internal understanding of intrinsic properties [bhattad2024stylegan, du2023generative, ke2024repurposing, ye2024stablenormal, anonymous2024intrinsiccontrolnet, luo2024intrinsicdiffusion]. We similarly leverage diffusion models’ ability to interpret intrinsic maps to generate realistic images as a foundation for rendering image composites.
| Guiding shadow | IC-Light [zhang2025iclight] | DiLightNet [zeng2024dilightnet] | Neural Gaffer [jin2024neural_gaffer] | ZeroComp+SDEdit | Ours | Ground truth |
@@ 3 Background: diffusion renderers
Recent work has shown that diffusion models can be adapted to generate photorealistic images from intrinsic maps including materials (e.g., albedo, roughness, metalness), geometry (e.g., surface normals and depth), and shading. This capability was built using a ControlNet in ZeroComp [zhang2025zerocomp] or full finetuning in RGBX [zeng2024rgb]. We refer to this new class of conditional generation models as “diffusion renderers”.
Specifically, ZeroComp [zhang2025zerocomp], which we will use in sec. 5, is trained to generate images from albedo, normals, depth, and partially-masked shading. It can be used for compositing in a zero-shot manner by alpha-compositing the intrinsics of the background and the object, and by masking out a region on and around the object in the shading map. The network then generates the full image with realistic shading. The intrinsics of the object are rendered using simple graphics shaders, while those of the background are estimated using pre-trained networks (e.g., [bhat2023zoedepth] for depth, see sec. 5.2). In the case of RGBX [zeng2024rgb], their XRGB network achieves object insertion by accepting as input alpha-composited albedo, normals, and a masked image. They then train an inpainting version of their network to generate the final image. Unlike traditional rendering, these methods offer no control over the lighting conditions.
In this work, we use such a pre-trained diffusion renderer and show that, by incorporating an approximate guiding shadow, we can steer the diffusion model to enable local lighting control, without any additional training. We now describe our approach in detail.
@@ 4 SpotLight
We employ a user-provided, guiding shadow to steer a pre-trained diffusion renderer (c.f. sec. 3) to follow the desired lighting direction. Our key insight is that even an approximate shadow encapsulates enough important lighting information and, by integrating this cue in the denoising process, we can simultaneously refine the composite appearance and its shadow, yielding a natural blend with the scene.
@@ 4.1 Approach overview
An overview of SpotLight is illustrated in fig. 2. In addition to the intrinsic maps for the backbone diffusion renderer, our approach requires two inputs: an object mask and a user-specified guiding shadow which guides the latent diffusion process. The guiding shadow can be obtained in several ways—in this paper, we experiment with fast rasterization [eisemann2011realtime], a trained shadow generation model [tasar2024controllable], and hand-drawn scribbles. Without retraining the diffusion renderer, we steer its generative process by incorporating the guiding shadow into the latent space.
4.2 Blending shadows with the background
Provided the guiding shadow , we first need to ensure that the diffusion model will properly blend the shadow with the background. We take inspiration from Blended Latent Diffusion [avrahami2023blendedlatent] to progressively blend the noisy latents at time with a noised VAE-encoded image which contains the shadow roughly composited over the image, . In practice, we obtain by compositing the object albedo and the guiding shadow on the background. The update made at each timestep to incorporate the desired shadow latents is
where is the shadow latent weight and is the shadow mask, bilinearly downsampled to the latent resolution (), with the edge opacity increased and dilated to maintain the desired shadow softness (see supp. for details). Noise is added to the encoded composite using the DDIM [song2020denoising] noise scheduler. Different from [avrahami2023blendedlatent], we do not use the slow test-time decoder optimization nor the progressive mask shrinking steps.
4.3 Enhancing lighting control
Simply running a diffusion renderer conditioned on the desired shadow direction yields perceptible, but subtle visual changes in the generated object (as will be shown in sec. 5.6). We propose to amplify the effect by running two parallel branches of the model: one which receives the shadow in the desired direction (positive branch), while the other has the shadow in the opposite direction (negative branch), obtained by shifting the light azimuth angle by . Although the opposite light direction works best for the negative branch, we found that casting no shadow also provides a good negative sample. From the outputs of the diffusion renderer (typically, -predictions [salimans2022progressive], i.e., linear combination of noise and estimated image), we then use classifier-free guidance to amplify the effect of the positive shadow on the lighting on the object:
where is the guidance scale, is the object mask bilinearly downsampled to the latent resolution (). After eq. 2, we run a regular DDIM sampling step to obtain the latents .
@@ 4.4 Final image synthesis
After the guided diffusion process, the resulting latents are decoded with the VAE decoder. The background preservation strategy of [zhang2025zerocomp] is applied to ensure that only the object and its shadow are modified, yielding a composite image with natural local lighting.
@@ 5 Evaluation on 3D object compositing
Evaluating composition methods where lighting conditions can be controlled is challenging, as the quality of the results must be assessed across multiple lighting condition for the same scene. To address this, we build over the evaluation setup of [zhang2025zerocomp] where 3D objects are realistically rendered and composited into background images using a physics-based renderer and ground-truth lighting extracted from HDR panoramas. We evaluate two scenarios: (1) “reference-based”, which preserves the original scene lighting to assess the ability of methods to produce images close to the ground truth, and (2) “user-controlled”, where the original lighting is modified (e.g., dominant light rotation) to simulate user-specified lighting conditions, requiring methods to adapt while maintaining a perceptually pleasing render. In this section, we describe how these scenarios are used for quantitative evaluation and a user study to assess the realism and controllability of the methods.
@@ 5.1 Evaluation dataset
We create datasets for the two scenarios using the Blender scenes graciously provided by the authors of [zhang2025zerocomp]. We first improve the 3D object rendered shadow by replacing the planar shadow catcher by a scene mesh using estimated depth from Depth Anything V2 [depth_anything_v2]. We also improve the ground truth intrinsic maps with antialiased versions.
“Reference-based” scenario. To assess the capacity of the methods to generate a composite in the ground truth lighting conditions, this scenario uses renders with the depth-warped HDR as in [zhang2025zerocomp] to ensure lighting fidelity around the inserted object. We extract the dominant light direction and use it to render the guiding shadow using blender’s EEVEE rasterization-based rendering engine, which employs real-time shadow mapping [eisemann2011realtime]. A total of 210 images featuring various objects are rendered, with example samples shown in fig. 3, in the “ground truth” column. This dataset is intended for quantitative evaluation (sec. 5.4).
“User-controlled” scenario. This dataset simulates a scenario where a user specifies a desired light position. 8 lighting directions are defined, spaced at increments in azimuth around the object, and used to create the guiding shadows, again using blender’s EEVEE engine. The same set of background and object combinations as in the “reference-based” version is used, resulting in a total of 1,680 images. This dataset is used in the user study (sec. 5.5) to evaluate the perceptual accuracy of renders under modified lighting conditions.